OpenStudy (hellomynameiskam):

Which ratio is not equivalent to 45/120 ? A.3/8 B. 5/12 C.9/24 D.90/240

OpenStudy (swamijatin2):



OpenStudy (anonymous):

what do u think???

OpenStudy (anonymous):

A its not correct

OpenStudy (hellomynameiskam):

i think its C ?

OpenStudy (swamijatin2):

9/24 =3/8

OpenStudy (inowalst):

> Divide the percent by 100 to convert it to a decimal. For example, if you want to find 20 percent of 5/7, first divide 20 by 100 to get 0.2. > Multiply the decimal by the numerator of the fraction. In this example, multiply 0.2 by 5 to get 1. > Place the result from the previous step over the original denominator. In this example, you would place 1 over 7 to find that 1/7 equals 20 percent of 5/7. If you want to convert to a decimal, simply divide the numerator by the denominator. In this example, 5/7 equals 0.7143. http://www.ehow.com/how_8119940_percent-fraction.html

OpenStudy (mathteacher1729):

How confident do you feel about reducing fractions? If you can reduce fractions, then this problem basically says: "reduce all the fractions you see, which one is equivalent to 45/120 (reduced)?"

OpenStudy (inowalst):

Oh and next time, no direct answers.^^

OpenStudy (hellomynameiskam):

D? I'm not good at reducing.

OpenStudy (mathteacher1729):

Do you know how to factor numbers? Like, 45 --> 45 = 9*5 45 = 3*3*5 That's it. You can't factor it anymore. And for 120 --> 120 = 2*60 120 = 2*2*30 120 = 2*2*2*15 120 = 2*2*2*3*5 Woooah. 45 and 120 have some factors in common -- 3 and 5. \[ \huge \dfrac{45}{120}=\frac{3*\color{red}{3*5}}{2*2*2*\color{red}{3*5}} \] Cancel the 3*5 in the numerator and denominator \[ \huge \dfrac{45}{120}=\frac{3}{2*2*2} = \frac{3}{8} \] Does that help?

OpenStudy (hellomynameiskam):

Yeah. so it's A

OpenStudy (mathteacher1729):

Nope. The problem says: > Which ratio is NOT equivalent to 45/120 ?

OpenStudy (swamijatin2):

so what i have said but no one think about my answer

OpenStudy (mathteacher1729):

And we just showed that 45/120 is equal to 3/8 , so those fractions ARE equivalent.

OpenStudy (mathteacher1729):

So the problem is asking this: "Reduce all the fractions. Which one does NOT reduce to 3/8?"

OpenStudy (hellomynameiskam):

B?

OpenStudy (swamijatin2):

yes B

OpenStudy (hellomynameiskam):

Kay

OpenStudy (mathteacher1729):

Did you reduce all the other fractions? (B Is correct, but I want to make sure you understand why it is correct.)